# thanks: Corresponding author

### v3

@@ III.5.3 Numerical results
$2$ $[-35.30,13.12]$ $[-29.03,11.28]$ $[-8.54,3.67]$ $[-5.38,3.01]$
$f_{T_{0}}(f_{T_{1}})/\Lambda^{4}$ $3$ $[-39.79,17.62]$ $[-32.85,15.10]$ $[-9.76,4.89]$ $[-6.31,3.95]$
$5$ $[-47.31,25.14]$ $[-39.22,21.47]$ $[-11.79,6.92]$ $[-7.90,5.53]$
$2$ $[-90.53,21.33]$ $[-75.97,17.78]$ $[-21.26,6.21]$ $[-12.00,5.22]$
$f_{T_{2}}/\Lambda^{4}$ $3$ $[-98.78,29.59]$ $[-82.88,24.69]$ $[-23.57,8.53]$ $[-13.78,7.00]$
$5$ $[-113.06,43.87]$ $[-94.83,36.64]$ $[-27.52,12.48]$ $[-16.83,10.04]$
$2$ $[-9.04,2.44]$ $[-7.81,2.07]$ $[-2.11,0.70]$ $[-1.26,0.58]$
$f_{T_{5}}(f_{T_{6}})/\Lambda^{4}$ $3$ $[-9.95,3.35]$ $[-8.59,2.84]$ $[-2.36,0.96]$ $[-1.45,0.78]$
$5$ $[-11.51,4.92]$ $[-9.93,4.18]$ $[-2.79,1.38]$ $[-1.78,1.11]$
$2$ $[-23.68,3.76]$ $[-16.59,3.08]$ $[-5.08,1.10]$ $[-2.84,0.97]$
$f_{T_{7}}/\Lambda^{4}$ $3$ $[-25.26,5.34]$ $[-17.85,4.33]$ $[-5.52,1.53]$ $[-3.20,1.33]$
$5$ $[-28.07,8.15]$ $[-20.06,6.54]$ $[-6.28,2.30]$ $[-3.82,1.95]$
$2$ $[-1.46,0.39]$ $[-1.25,0.33]$ $[-0.33,0.11]$ $[-0.20,0.09]$
$f_{T_{8}}/\Lambda^{4}$ $3$ $[-1.61,0.54]$ $[-1.38,0.46]$ $[-0.37,0.15]$ $[-0.23,0.12]$
$5$ $[-1.86,0.79]$ $[-1.59,0.68]$ $[-0.44,0.22]$ $[-0.28,0.18]$
$2$ $[-3.92,0.60]$ $[-2.78,0.50]$ $[-0.94,0.18]$ $[-0.48,0.16]$
$f_{T_{9}}/\Lambda^{4}$ $3$ $[-4.17,0.85]$ $[-2.99,0.71]$ $[-1.01,0.25]$ $[-0.54,0.22]$
$5$ $[-4.63,1.31]$ $[-3.35,1.07]$ $[-1.14,0.38]$ $[-0.64,0.32]$
@@
$2$ $[-77.26,28.50]$ $[-37.51,13.86]$
$f_{T_{0}}(f_{T_{1}})/\Lambda^{4}$ $3$ $[-87.06,38.30]$ $[-42.27,18.63]$
$5$ $[-103.48,54.72]$ $[-50.26,26.61]$
$2$ $[-197.39,46.98]$ $[-90.17,22.29]$
$f_{T_{2}}/\Lambda^{4}$ $3$ $[-215.55,65.15]$ $[-98.72,30.84]$
$5$ $[-246.95,96.54]$ $[-113.45,45.58]$
$2$ $[-19.41,5.37]$ $[-9.28,2.53]$
$f_{T_{5}}(f_{T_{6}})/\Lambda^{4}$ $3$ $[-21.41,7.37]$ $[-10.23,3.48]$
$5$ $[-24.83,10.79]$ $[-11.85,5.11]$
$2$ $[-47.96,8.15]$ $[-22.65,3.91]$
$f_{T_{7}}/\Lambda^{4}$ $3$ $[-51.34,11.52]$ $[-24.27,5.52]$
$5$ $[-57.33,17.52]$ $[-27.13,8.39]$
$2$ $[-3.08,0.84]$ $[-1.48,0.40]$
$f_{T_{8}}/\Lambda^{4}$ $3$ $[-3.39,1.15]$ $[-1.63,0.55]$
$5$ $[-3.93,1.69]$ $[-1.89,0.81]$
$2$ $[-8.92,1.33]$ $[-3.86,0.63]$
$f_{T_{9}}/\Lambda^{4}$ $3$ $[-9.49,1.89]$ $[-4.13,0.89]$
$5$ $[-10.50,2.90]$ $[-4.60,1.36]$
@@
Using the designed luminosities for both the “conservative” and “optimistic” cases [125, 126] and the fitted cross-sections, the expected constraints can be estimated using Eq. (4). The results are listed in Tables 12 and 13. It can be shown that the VQSN is effective in the search for aQGCs, the expected constraints generally outperform the ones obtained by a traditional event selection strategy [151].
@@
IV Summary
@@
In this study, we present a novel (V)QSN algorithm for classification that is designed to operate on a quantum computer. A numerical example running on the tianyan176-2 of the Tianyan quantum computing cloud platform is established to show that the VQSN is suitable for a general proposed classification. This algorithm is analogous to the KNN algorithm. We employ the (V)QSN and the (t-)KNN with the objective of identifying signals of NP in different processes.
@@
The VQSN is used in the phenomenological study of the gQGCs in the process $\mu^{+}\mu^{-}\to jj\nu\bar{\nu}$ at a $\sqrt{s}=3$ TeV muon collider. The VQSN event selection strategy is tested on the tianyan176-2. The expected coefficient constraints are obtained and found to be at the same order of magnitude as the traditional event selection strategy in Ref. [124], indicating that the VQSN is an effective method and possesses noise tolerance.
@@
The gQGCs at the LHC are also studied, using the processes $pp\to jj\ell^{+}\ell^{-}\nu\bar{\nu}$, $pp\to j\ell^{+}\ell^{-}\nu\bar{\nu}$ and $pp\to\gamma j\ell^{+}\ell^{-}\nu\bar{\nu}$. These processes can be concluded to be sensitive to the gQGCs. The constraints on the coefficient are of a similar magnitude to those obtained in the $pp\to\gamma\gamma$ process [91]. Consequently, they can serve as a valuable supplement to the constraints on the gQGCs. Moreover, the $pp\to jj\ell^{+}\ell^{-}\nu\bar{\nu}$ process contains a VBS process which is expected to be more significant at higher energies [151], therefore is expected to be more important at future colliders such as the HE-LHC and FCC-hh [154, 155].
@@
Another application is the contribution of the aQGCs in the tri-photon processes at muon colliders. In this case, only the VQSN is studied. The expected constraints are obtained, and found to be more strengthen than the ones obtained using an traditional event selection strategy [151].
@@
In terms of cc, the VQSN is capable of achieving the same order of magnitude of the discriminative power to that of the KNN, while requiring one order of magnitude less cc in the three applications. It is important to note that the datasets used in this study are classical, i.e., even when working with classical data, the VQSN can demonstrate its superiority in efficiency. When keeping the amount of test data much larger than the amount of training data, the larger training dataset and the larger feature space, the more efficient the VQSN is, the significance of VQSN is likely to increase as the processes under consideration become more intricate.
@@
This study has not addressed further improvements in encoding the datasets in optimistic ways [86, 87] which are suitable for cooperation with (V)QSN. Furthermore, it would be beneficial to investigate the simulation of the collision process on a quantum computer. Reading results directly from a quantum computer is resource intensive, QSN can process data within the quantum computer, avoiding this consumption. In particular, it is plausible that the training dataset originates from simulations on a quantum computer, while the test dataset is derived from classical data obtained from a collider. This scenario is well-suited to the QSN.
@@ Acknowledgements.
We sincerely thank the Tianyan quantum computing cloud platform of China Telecom Quantum Information Technology Group Co., LTD. for providing quantum computer machine time and technical support. This work was supported in part by the National Natural Science Foundation of China under Grants Nos. 11875157 and 12147214, and the Natural Science Foundation of the Liaoning Scientific Committee No. LJKZ0978.

### v4

@@ 3.5.3 Numerical results
22 [35.30,13.12][-35.30,13.12] [29.03,11.28][-29.03,11.28] [8.54,3.67][-8.54,3.67] [5.38,3.01][-5.38,3.01]
$f_{T_{0}}(f_{T_{1}})/\Lambda^{4}$ 33 [39.79,17.62][-39.79,17.62] [32.85,15.10][-32.85,15.10] [9.76,4.89][-9.76,4.89] [6.31,3.95][-6.31,3.95]
55 [47.31,25.14][-47.31,25.14] [39.22,21.47][-39.22,21.47] [11.79,6.92][-11.79,6.92] [7.90,5.53][-7.90,5.53]
22 [90.53,21.33][-90.53,21.33] [75.97,17.78][-75.97,17.78] [21.26,6.21][-21.26,6.21] [12.00,5.22][-12.00,5.22]
$f_{T_{2}}/\Lambda^{4}$ 33 [98.78,29.59][-98.78,29.59] [82.88,24.69][-82.88,24.69] [23.57,8.53][-23.57,8.53] [13.78,7.00][-13.78,7.00]
55 [113.06,43.87][-113.06,43.87] [94.83,36.64][-94.83,36.64] [27.52,12.48][-27.52,12.48] [16.83,10.04][-16.83,10.04]
22 [9.04,2.44][-9.04,2.44] [7.81,2.07][-7.81,2.07] [2.11,0.70][-2.11,0.70] [1.26,0.58][-1.26,0.58]
$f_{T_{5}}(f_{T_{6}})/\Lambda^{4}$ 33 [9.95,3.35][-9.95,3.35] [8.59,2.84][-8.59,2.84] [2.36,0.96][-2.36,0.96] [1.45,0.78][-1.45,0.78]
55 [11.51,4.92][-11.51,4.92] [9.93,4.18][-9.93,4.18] [2.79,1.38][-2.79,1.38] [1.78,1.11][-1.78,1.11]
22 [23.68,3.76][-23.68,3.76] [16.59,3.08][-16.59,3.08] [5.08,1.10][-5.08,1.10] [2.84,0.97][-2.84,0.97]
$f_{T_{7}}/\Lambda^{4}$ 33 [25.26,5.34][-25.26,5.34] [17.85,4.33][-17.85,4.33] [5.52,1.53][-5.52,1.53] [3.20,1.33][-3.20,1.33]
55 [28.07,8.15][-28.07,8.15] [20.06,6.54][-20.06,6.54] [6.28,2.30][-6.28,2.30] [3.82,1.95][-3.82,1.95]
22 [1.46,0.39][-1.46,0.39] [1.25,0.33][-1.25,0.33] [0.33,0.11][-0.33,0.11] [0.20,0.09][-0.20,0.09]
$f_{T_{8}}/\Lambda^{4}$ 33 [1.61,0.54][-1.61,0.54] [1.38,0.46][-1.38,0.46] [0.37,0.15][-0.37,0.15] [0.23,0.12][-0.23,0.12]
55 [1.86,0.79][-1.86,0.79] [1.59,0.68][-1.59,0.68] [0.44,0.22][-0.44,0.22] [0.28,0.18][-0.28,0.18]
22 [3.92,0.60][-3.92,0.60] [2.78,0.50][-2.78,0.50] [0.94,0.18][-0.94,0.18] [0.48,0.16][-0.48,0.16]
$f_{T_{9}}/\Lambda^{4}$ 33 [4.17,0.85][-4.17,0.85] [2.99,0.71][-2.99,0.71] [1.01,0.25][-1.01,0.25] [0.54,0.22][-0.54,0.22]
55 [4.63,1.31][-4.63,1.31] [3.35,1.07][-3.35,1.07] [1.14,0.38][-1.14,0.38] [0.64,0.32][-0.64,0.32]
@@
22 [77.26,28.50][-77.26,28.50] [37.51,13.86][-37.51,13.86]
$f_{T_{0}}(f_{T_{1}})/\Lambda^{4}$ 33 [87.06,38.30][-87.06,38.30] [42.27,18.63][-42.27,18.63]
55 [103.48,54.72][-103.48,54.72] [50.26,26.61][-50.26,26.61]
22 [197.39,46.98][-197.39,46.98] [90.17,22.29][-90.17,22.29]
$f_{T_{2}}/\Lambda^{4}$ 33 [215.55,65.15][-215.55,65.15] [98.72,30.84][-98.72,30.84]
55 [246.95,96.54][-246.95,96.54] [113.45,45.58][-113.45,45.58]
22 [19.41,5.37][-19.41,5.37] [9.28,2.53][-9.28,2.53]
$f_{T_{5}}(f_{T_{6}})/\Lambda^{4}$ 33 [21.41,7.37][-21.41,7.37] [10.23,3.48][-10.23,3.48]
55 [24.83,10.79][-24.83,10.79] [11.85,5.11][-11.85,5.11]
22 [47.96,8.15][-47.96,8.15] [22.65,3.91][-22.65,3.91]
$f_{T_{7}}/\Lambda^{4}$ 33 [51.34,11.52][-51.34,11.52] [24.27,5.52][-24.27,5.52]
55 [57.33,17.52][-57.33,17.52] [27.13,8.39][-27.13,8.39]
22 [3.08,0.84][-3.08,0.84] [1.48,0.40][-1.48,0.40]
$f_{T_{8}}/\Lambda^{4}$ 33 [3.39,1.15][-3.39,1.15] [1.63,0.55][-1.63,0.55]
55 [3.93,1.69][-3.93,1.69] [1.89,0.81][-1.89,0.81]
22 [8.92,1.33][-8.92,1.33] [3.86,0.63][-3.86,0.63]
$f_{T_{9}}/\Lambda^{4}$ 33 [9.49,1.89][-9.49,1.89] [4.13,0.89][-4.13,0.89]
55 [10.50,2.90][-10.50,2.90] [4.60,1.36][-4.60,1.36]
@@
Using the designed luminosities for both the “conservative” and “optimistic” cases Black:2022cth ; Accettura:2023ked and the fitted cross-sections, the expected constraints can be estimated using Eq. (4). The results are listed in Tables 12 and 13. It can be shown that the VQSN is effective in the search for aQGCs, the expected constraints generally outperform the ones obtained by a traditional event selection strategy Yang:2020rjt .
@@
3.6 Discussion on efficacy difference between (t-)KNN and (V)QSN
In studies of several new physics scenarios, numerical results indicate that t-KNN achieves tighter constraints on coefficients compared to VQSN. Through testing, we identified that such difference comes from the fundamental distinction: t-KNN is intrinsically an algorithm considering only nearest neighbors in training dataset, whereas VQSN employs a weighted scoring over the entire training dataset.
While the VQSN method can technically be adapted to classify using only nearest neighbors basheer2021quantumknearestneighborsalgorithm , implementing this quantum variant demands additional gate operations. This not only fails to surpass the computational efficiency of classical t-KNN when processing classical data but also produces deeper circuits impractical for execution on current noisy quantum hardware. Consequently, we contend that a quantum algorithm achieving coefficient constraints on par with t-KNN (within the same order of magnitude) while simultaneously reducing cc by an order of magnitude and remaining executable on current quantum hardware would hold significant value for high-luminosity collider experiments confronting exponentially growing datasets.
Figure 18: Two examples that global weighted can surpass the KNN. The left panels are the points in training dataset with the labels $class_{i}=\pm 1$. The middle panels are the points in test dataset and decision boundary calculated by the sign of $\sum_{i\in training}class_{i}/\left(v_{test}-v_{i}\right)^{2}$. The right panels are the points in test dataset and decision boundary calculated by using KNN with $k=5$. For the test set, blue points within red regions represent misclassifications, and vice versa. Accuracy (denoted as Acc) is defined as the proportion of correctly classified data points in the test set.
Furthermore, there is no inherent superiority between algorithms that consider only nearest neighbors and those utilizing weighted scoring across the entire training dataset. The relative efficacy depends fundamentally on the data distribution of the problem under study. To facilitate discussion, we restrict our analysis to classical algorithms: one employing distance inverse weighted scoring and another using nearest neighbor based unweighted voting (KNN). In Fig. 18 we present two illustrative cases. For the upper row, points of both classes interpenetrate extensively. Within overlapping regions, some points are entirely surrounded by instances of the opposite class, causing KNN to perform worse than the global weighting approach. The bottom row in Fig. 18 relates less to distribution characteristics and more to training dataset sampling. Consider the lowermost point in the training data, due to accidentally sparse sampling in its neighborhood, the locally focused KNN underperforms compared to the globally-aware weighting method.
Therefore, given VQSN’s superior performance in cc and its potential to surpass t-KNN under specific conditions in efficacy, coupled with unexplored opportunities for feature space optimization, we conclude that VQSN represents a valuable algorithm for NP searches.
4 Summary
@@
In this study, we present a (V)QSN algorithm for classification that is designed to operate on a quantum computer. A numerical example running on the tianyan176-2 of the Tianyan quantum computing cloud platform is established to show that the VQSN is suitable for a general proposed classification. This algorithm is analogous to the KNN algorithm. We employ the (V)QSN and the (t-)KNN with the objective of identifying signals of NP in different processes.
@@
The VQSN is used in the phenomenological study of the gQGCs in the process $\mu^{+}\mu^{-}\to jj\nu\bar{\nu}$ at a $\sqrt{s}=3$ TeV muon collider. The VQSN event selection strategy is tested on the tianyan176-2. The expected coefficient constraints are obtained and found to be at the same order of magnitude as the traditional event selection strategy in Ref. Yang:2023gos , indicating that the VQSN is an effective method and possesses noise tolerance.
@@
The gQGCs at the LHC are also studied, using the processes $p\to jj\ell^{+}\ell^{-}\nu\bar{\nu}, pp$$p\to j\ell^{+}\ell^{-}\nu\bar{\nu}$ and $pp\to\gamma j\ell^{+}\ell^{-}\nu\bar{\nu}$. These processes can be concluded to be sensitive to the gQGCs. The constraints on the coefficient are of a similar magnitude to those obtained in the $pp\to\gamma\gamma$ process Ellis:2018cos . Consequently, they can serve as a valuable supplement to the constraints on the gQGCs. Moreover, the $pp\to jj\ell^{+}\ell^{-}\nu\bar{\nu}$ process contains a VBS process which is expected to be more significant at higher energies Yang:2020rjt , therefore is expected to be more important at future colliders such as the HE-LHC and FCC-hh FCC:2018bvk ; Benedikt:2022kan .
@@
Another application is the contribution of the aQGCs in the tri-photon processes at muon colliders. In this case, only the VQSN is studied. The expected constraints are obtained, and found to be more strengthen than the ones obtained using an traditional event selection strategy Yang:2020rjt .
@@
In terms of cc, the VQSN is capable of achieving the same order of magnitude of the discriminative power to that of the KNN, while requiring one order of magnitude less cc in the three applications. It is important to note that the datasets used in this study are classical, i.e., even when working with classical data, the VQSN can demonstrate its superiority in efficiency in the sense of cc. When keeping the amount of test data much larger than the amount of training data, the larger training dataset and the larger feature space, the more efficient the VQSN is, the significance of VQSN is likely to increase as the processes under consideration become more intricate.
@@
This study has not addressed further improvements in encoding the datasets in optimistic ways Havlicek:2018nqz ; Lloyd:2020eeh which are suitable for cooperation with (V)QSN. Furthermore, it would be beneficial to investigate the simulation of the collision process on a quantum computer. Reading results directly from a quantum computer is resource intensive, QSN can process data within the quantum computer, avoiding this consumption. In particular, it is plausible that the training dataset originates from simulations on a quantum computer, while the test dataset is derived from classical data obtained from a collider. This scenario is well-suited to the QSN.
@@ Acknowledgements.
We sincerely thank the Tianyan quantum computing cloud platform of China Telecom Quantum Information Technology Group Co., LTD. for providing quantum computer machine time and technical support. This work was supported in part by the National Natural Science Foundation of China under Grants Nos. 12575106, 11875157 and 12147214, and the Natural Science Foundation of the Liaoning Scientific Committee No. LJKZ0978.